Prepare 50ml of 0.3N Ba(OH)2 * 8 H2O


1
Expert's answer
2021-09-14T01:13:02-0400

Normality = "\frac{no.of.equivalent×1000}{volume.of.solution(mL)}"


0.3 = "\frac{no.of.equivalent×1000}{50}"


no. of equivalent = 0.015


no. of equivalent = no. of moles/ n factor

0.015 = no. of moles / 2

no. of moles = 0.030


( Gram required / molar mass )= 0.030

Gram required = 0.030 × 315.46

Gram required = 9.46 g


9.46 g of Ba(OH)2 is required to prepare 0.3N solution